# Analog model of Kerr black holes in scalar-tensor theories using binary Bose-Einstein condensates: Quasibound states and tachyonic instabilities

### v1

@@ Acknowledgements.
$h_{H}=\bigg{\{}\begin{array}[]{ll}e^{-i(\omega-m\ell)r^{\ast}},&r^{\ast}\rightarrow-\infty,\\ A_{\infty}^{-}e^{-i\omega r^{\ast}}+A_{\infty}^{+}e^{i\omega r^{\ast}},&r^{\ast}\rightarrow\infty,\end{array}$ (S19)
@@
$h_{\infty}=\bigg{\{}\begin{array}[]{ll}A_{H}^{-}e^{-i(\omega-m\ell)r^{\ast}}+A_{H}^{+}e^{i(\omega-m\ell)r^{\ast}},&r^{\ast}\rightarrow-\infty,\\ e^{i\omega r^{\ast}},&r^{\ast}\rightarrow\infty,\end{array}$ (S22)
respectively. The spectrum then can be obtained from the condition $A_{H}^{+}/A_{H}^{-}=0$ [44, 45]. According to the effective mass term we chose in (6), it is suitable to adopt the generalized Pöschl-Teller potential by splitting a whole space into three regions through the following parametrization [45]
$V_{\text{PT}}(r^{\ast})=$ $V_{3}(r^{\ast})\Theta(\bar{r}^{\ast}-r^{\ast})+V_{2}(r^{\ast})[\Theta(r^{\ast}-\bar{r}^{\ast})-\Theta(r^{\ast}-\tilde{r}^{\ast})]$
$+V_{1}(r^{\ast})\Theta(r^{\ast}-\tilde{r}^{\ast})\,,$ (S23)
@@
$V_{i}(r^{\ast})=a_{i}+b_{i}\text{sech}^{2}{[(r^{\ast}-{r}_{i}^{\ast})/s_{i}]}\,$ (S24)
with $r_{1}^{\ast}=r_{0}^{\ast},r_{2}^{\ast}=\tilde{r}^{\ast},r_{3}^{\ast}=\bar{r}^{\ast}$. $\bar{r}^{\ast}$ is the position of the top of the rotational barrier, and $\tilde{r}^{\ast}$ is selected to be within the interval $\{\bar{r}^{\ast},r_{0}^{\ast}\}$ where $r_{0}^{\ast}$ is the position of another barrier induced by the mass term in Fig. 4. The parameters can be determined by comparing with the effective potential $V_{\text{eff}}$ with the mass term in (6) in the text. The coefficients $a_{1}$ and $b_{1}$ can also be determined by $r^{\ast}\rightarrow\infty$ and the mass shell formula giving $a_{1}=\omega^{2}$ and $b_{1}=-2\Omega_{0}$. Likewise, the coefficients $a_{1}$ and $b_{1}$ are obtained from the effective potential near the horizon as well as at the top of the rotational barrier giving $a_{3}=(\omega-ml)^{2}$ and $b_{3}=V_{\text{eff}}(\bar{r}^{\ast})-a_{3}$. Finally, considering the barrier given by the mass shell is far from the rotational barrier, $a_{2}=\omega^{2}$ and the coefficient $b_{2}$ can be related to the potential value at the rotational barrier by $b_{2}=V_{\text{eff}}(\bar{r}^{\ast})-a_{2}$. The width $s_{3}$ is the mass shell width $1/\alpha$. The widths $s_{1,2}$ are chosen to make $V_{\text{PT}}$ is differentiable at the top of barrier $\bar{r}^{\ast}$ [45]. In each regions, the general solution to the Schrodinger-like equation
$\partial_{r^{\ast}}^{2}H(r^{\ast})+V_{i}H(r^{\ast})=0$ (S25)
is given by the linear combination of two independent solutions [46]
$H_{i}(r^{\ast})=A_{i}h_{i}^{\text{in}}(r^{\ast})+B_{i}h_{i}^{\text{up}}(r^{\ast}),$ (S26)
@@
$\beta_{i}=-\frac{1}{2}+\frac{\sqrt{1+4s_{i}^{2}b_{i}}}{2}.$ (S27)
Given the condition of (S22) at $r^{\ast}\rightarrow\infty$, it requires $A_{1}=0,B_{1}=1$. We then match the functions $h_{1}^{\text{(up)}}$, $h_{2}^{\text{(in)}}$, $h_{2}^{\text{(up)}}$, $h_{3}^{\text{(in)}}$, and $h_{3}^{\text{(out)}}$ at two matching points: one is the top of the rotational barrier $\bar{r}^{\ast}$ and the other is $\tilde{r}^{\ast}$, say at $\tilde{r}^{\ast}=({r}_{0}^{\ast}+\tilde{r}^{\ast})/2$ given by
$h_{1}^{\text{(up)}}\big{|}_{\tilde{r}^{\ast}}={A}_{2}h_{2}^{\text{(in)}}\big{|}_{\tilde{r}^{\ast}}+{B}_{2}h_{2}^{\text{(up)}}\big{|}_{\tilde{r}^{\ast}},$ (S28)
$\partial_{r^{\ast}}h_{1}^{\text{(up)}}\big{|}_{\tilde{r}^{\ast}}={A}_{2}\partial_{{r}^{\ast}}h_{2}^{\text{(in)}}|_{\tilde{r}^{\ast}}+{B}_{2}\partial_{{r}^{\ast}}h_{2}^{\text{(up)}}\big{|}_{\tilde{r}^{\ast}},$ (S29)
${A}_{3}h_{3}^{\text{(in)}}+{B}_{3}h_{3}^{\text{(up)}}\big{|}_{\bar{r}^{\ast}}=\mathcal{A}_{2}h_{2}^{\text{(in)}}\big{|}_{\bar{r}^{\ast}}+{B}_{2}h_{2}^{\text{(up)}}\big{|}_{\bar{r}^{\ast}},$ (S30)
${A}_{3}\partial_{r^{\ast}}h_{3}^{\text{(in)}}\big{|}_{\bar{r}^{\ast}}+{B}_{3}\partial_{r^{\ast}}h_{3}^{\text{(up)}}\big{|}_{\bar{r}^{\ast}}$ (S31)
@@
Now the coefficients $A_{3}$ and $B_{3}$ can be determined from above matching conditions. To read out $A_{H}^{+}$ and $A_{H}^{-}$ for the asymptotical behavior of the solution at $r^{\ast}\rightarrow-\infty$ in (S22) from the solution at $i=3$ in (S26), one more step is to rewrite $h_{3}^{\text{(up)}}$ into a combination of the following plane wave forms by employing the identity of the hypergeometric functions
${}_{2}F_{1}(a,b;c;y)=$ $\frac{\Gamma(c)\Gamma(c-a-b)}{\Gamma(c-a)\Gamma(c-b)}$
×$\times_{2}F_{1}(a,b;a+b+1-c;1-y)$
$+\frac{\Gamma(c)\Gamma(a+b-c)}{\Gamma(a)\Gamma(b)}(1-y)^{c-a-b}$
×$\times_{2}F_{1}(c-a,c-b;1+c-a-b;1-y),$ (S32)
@@
×$\times_{2}F_{1}\left(-\beta,\beta+1;1-\mu;\frac{1-y}{2}\right),$ (S33)
@@
$h_{3}^{\text{(}up)}(r^{\ast}\rightarrow-\infty)={N}e^{-\mu_{3}r^{\ast}/s_{3}}+{M}e^{\mu_{3}r^{\ast}/s_{3}}$ (S34)
@@
${M}=\frac{\Gamma(1-\mu_{3})\Gamma(-\mu_{3})}{\Gamma(1-\mu_{3}+\beta_{3})\Gamma(-\mu_{3}-\beta_{3})}.$ (S35)
@@
$A_{H}^{-}={A}_{3}+{N}{B}_{3},$ (S36)
$A_{H}^{+}={M}+{B}_{3}.$ (S37)
When the mass shell is absent, $A_{3}$ and $B_{3}$ are found in the matching conditions with ${A}_{2}=0$, ${B}_{2}=1$. The spectrum then can be found by $A_{H}^{+}/A_{H}^{-}=0$ [44, 45]. Alternatively, one can start from the $i=3$ region and use (S19) to determine $A_{3}=1$ and $B_{3}=0$. The matching points are still at $\bar{r}^{\ast}$ and $\tilde{r}^{\ast}$. Using the above identity rewrites $h_{1}^{\text{(}in)}$ at $r^{\ast}\rightarrow\infty$. Then $A^{(+)}_{\infty}$ and $A^{(-)}_{\infty}$ can be read off from (S19). The requirement of $A^{(-)}_{\infty}/A^{(+)}_{\infty}=0$ gives the eigenfequency.
Figure 5: The frequencies of the quasinormal mode of the fundamental tone for various uniform Rabi coupling strengths, and $m=1,2,3$.
With reference to [42], we introduce another method suited for a constant mass term. Defining $x=r^{-1}$, we look for a solution
$H(x)=e^{i\tilde{{\omega}}x^{-1}}\left(\frac{1-x}{1+x}\right)^{-i\left(\omega-m\ell\right)/2}\sum_{k=0}^{\infty}a_{k}\left(1-x\right)^{k},$ (S38)
where $\tilde{\omega}=\sqrt{\omega^{2}-4\Omega_{0}^{2}}$, $a_{0}=1$ and $\Omega_{0}$ of the Rabi coupling strength is considered as a constant. After substituting it into the Schrodinger-like wave equation in the text, we have four-term recurrence relation given by
$\alpha_{0}a_{1}+\beta_{0}a_{0}=0,$ (S39)
$\alpha_{1}a_{2}+\beta_{1}a_{1}+\gamma_{1}a_{0}=0,$ (S40)
and
$\alpha_{k}a_{k+1}+\beta_{k}a_{k}+\gamma_{k}a_{k-1}+\delta_{k}a_{k-2}=0,$ (S41)
@@
$\alpha_{k}=$ $-8\left(1+k\right)\left[(1+k)-i\left(\omega-m\ell\right)\right]\,,$ (S42)
$\beta_{k}=$ $4\big{[}5k^{2}+k\left(5+4im\ell-4i\omega-4i\tilde{\omega}\right)+2m\ell\left(i+\omega+\tilde{\omega}\right),$
$+\left(i+m+\omega+\tilde{\omega}\right)\left(-i+m-\omega-\tilde{\omega}\right)\big{]},$ (S43)
$\gamma_{k}=$ $-2\left(-3+8k^{2}+4im\ell k-4i\omega k-4i\tilde{\omega}k\right).$ (S44)
$\delta_{k}=$ $(1+2k)(-3+2k).$ (S45)
When the series expansion in (S38) converges to a finite value within $0\leq x\leq 1$, the solution of $H(x)$ satisfies the QNM boundary condition with the asymptotic forms of $e^{i\tilde{\omega}/x}$ at the horizon ($x\rightarrow 1$) and $(1-x)^{-i(\omega-m\ell)/2}e^{i\tilde{\omega}/x}$ at infinity ($x\rightarrow 0$). Furthermore, one can use Gaussian elimination to reduce (S41) to the three-term recurrence relation
$\alpha_{0}^{\prime}a_{1}+\beta_{0}^{\prime}a_{0}=0,$ (S46)
$\alpha_{k}^{\prime}a_{k+1}+\beta_{k}^{\prime}a_{k}+\gamma_{k}^{\prime}a_{k-1}=0,\quad k=1,2,\cdots.$ (S47)
@@
$\alpha_{k}^{\prime},\,\beta_{k}^{\prime}$ and $\gamma_{k}^{\prime}$ can be expressed in terms of $\alpha_{k},\,\beta_{k},\,\gamma_{k}$, and $\delta_{k}$, given by
$\alpha_{0}=\alpha_{0}^{\prime},\,\beta_{0}^{\prime}=\beta,\,\gamma_{0}^{\prime}=\gamma_{0}$ (S48)
$\alpha_{1}=\alpha_{1}^{\prime},\,\beta_{1}^{\prime}=\beta_{1},\,\gamma_{1}^{\prime}=\gamma_{1}$ (S49)
$\alpha_{k}^{\prime}=\alpha_{k},\,\beta_{k}^{\prime}=\beta_{k}-\alpha_{k-1}^{\prime}\delta_{k}/\gamma_{k-1}^{\prime},\quad{and}\quad\gamma_{k}^{\prime}=\gamma_{k}-\beta_{k-1}^{\prime}\delta_{k}/\gamma_{k-1}^{\prime},$ (S50)
which gives the QNM condition written as
$\beta_{0}^{\prime}-\frac{\alpha_{0}^{\prime}\gamma_{1}^{\prime}}{\beta_{1}^{\prime}-}\frac{\alpha_{1}^{\prime}\gamma_{2}^{\prime}}{\beta_{2}^{\prime}-}\frac{\alpha_{2}^{\prime}\gamma_{3}^{\prime}}{\beta_{3}^{\prime}-}\cdots=0.$ (S51)
The complex-frequency roots of the above condition corresponding to a different uniform Rabi coupling strength are shown in Fig. 5. To compare with Fig. 1, the modification of the fundamental mode in a constant $\Omega_{0}$ are not as drastic as in the spatial-dependent one from the case of $\Omega_{0}=0$.
Numerically solving the time-dependent equation in (5) for $H_{m}(r,t)$

### v2

@@ Appendix B PÖSCHL-TELLER METHOD AND CONTINUED FRACTION METHOD
$h_{H}=\bigg{\{}\begin{array}[]{ll}e^{-i(\omega-m\ell)r^{\ast}},&r^{\ast}\rightarrow-\infty\\ A_{\infty}^{-}e^{-i\omega r^{\ast}}+A_{\infty}^{+}e^{i\omega r^{\ast}},&r^{\ast}\rightarrow\infty\end{array},$ (32)
@@
$h_{\infty}=\bigg{\{}\begin{array}[]{ll}A_{H}^{-}e^{-i(\omega-m\ell)r^{\ast}}+A_{H}^{+}e^{i(\omega-m\ell)r^{\ast}},&r^{\ast}\rightarrow-\infty\\ e^{i\omega r^{\ast}},&r^{\ast}\rightarrow\infty\end{array},$ (35)
respectively. The spectrum then can be obtained from the condition $A_{H}^{+}/A_{H}^{-}=0$ Macedo et al. (2018); Torres et al. (2022). According to the effective mass term we chose in (9), it is suitable to adopt the generalized Pöschl-Teller potential by splitting a whole space into three regions through the following parametrization Torres et al. (2022)
$V_{\text{PT}}(r^{\ast})=$ $V_{3}(r^{\ast})\Theta(\bar{r}^{\ast}-r^{\ast})$
$+V_{2}(r^{\ast})[\Theta(r^{\ast}-\bar{r}^{\ast})-\Theta(r^{\ast}-\tilde{r}^{\ast})]$
$+V_{1}(r^{\ast})\Theta(r^{\ast}-\tilde{r}^{\ast})\,,$ (36)
@@
$V_{i}(r^{\ast})=a_{i}+b_{i}\text{sech}^{2}{[(r^{\ast}-{r}_{i}^{\ast})/s_{i}]}\,$ (37)
with $r_{1}^{\ast}=r_{0}^{\ast},r_{2}^{\ast}=\tilde{r}^{\ast},r_{3}^{\ast}=\bar{r}^{\ast}$. $\bar{r}^{\ast}$ is the position of the top of the rotational barrier, and $\tilde{r}^{\ast}$ is selected to be within the interval $\{\bar{r}^{\ast},r_{0}^{\ast}\}$ where $r_{0}^{\ast}$ is the position of another barrier induced by the mass term in Fig. 5. The parameters can be determined by comparing with the effective potential $V_{\text{eff}}$ with the mass term in (9) in the text. The coefficients $a_{1}$ and $b_{1}$ can also be determined by $r^{\ast}\rightarrow\infty$ and the mass shell formula giving $a_{1}=\omega^{2}$ and $b_{1}=-2\Omega_{0}$. Likewise, the coefficients $a_{1}$ and $b_{1}$ are obtained from the effective potential near the horizon as well as at the top of the rotational barrier giving $a_{3}=(\omega-ml)^{2}$ and $b_{3}=V_{\text{eff}}(\bar{r}^{\ast})-a_{3}$. Finally, considering the barrier given by the mass shell is far from the rotational barrier, $a_{2}=\omega^{2}$ and the coefficient $b_{2}$ can be related to the potential value at the rotational barrier by $b_{2}=V_{\text{eff}}(\bar{r}^{\ast})-a_{2}$. The width $s_{3}$ is the mass shell width $1/\alpha$. The widths $s_{1,2}$ are chosen to make $V_{\text{PT}}$ is differentiable at the top of barrier $\bar{r}^{\ast}$ Torres et al. (2022). In each regions, the general solution to the Schrodinger-like equation
$\partial_{r^{\ast}}^{2}H(r^{\ast})+V_{i}H(r^{\ast})=0$ (38)
is given by the linear combination of two independent solutions Casals et al. (2009)
$H_{i}(r^{\ast})=A_{i}h_{i}^{\text{in}}(r^{\ast})+B_{i}h_{i}^{\text{up}}(r^{\ast}),$ (39)
@@
$\beta_{i}=-\frac{1}{2}+\frac{\sqrt{1+4s_{i}^{2}b_{i}}}{2}.$ (40)
Given the condition of (35) at $r^{\ast}\rightarrow\infty$, it requires $A_{1}=0,B_{1}=1$. We then match the functions $h_{1}^{\text{(up)}}$, $h_{2}^{\text{(in)}}$, $h_{2}^{\text{(up)}}$, $h_{3}^{\text{(in)}}$ and $h_{3}^{\text{(out)}}$ at two matching points: one is the top of the rotational barrier $\bar{r}^{\ast}$ and the other is $\tilde{r}^{\ast}$, say at $\tilde{r}^{\ast}=({r}_{0}^{\ast}+\tilde{r}^{\ast})/2$ given by
$h_{1}^{\text{(up)}}\big{|}_{\tilde{r}^{\ast}}={A}_{2}h_{2}^{\text{(in)}}\big{|}_{\tilde{r}^{\ast}}+{B}_{2}h_{2}^{\text{(up)}}\big{|}_{\tilde{r}^{\ast}},$ (41)
$\partial_{r^{\ast}}h_{1}^{\text{(up)}}\big{|}_{\tilde{r}^{\ast}}={A}_{2}\partial_{{r}^{\ast}}h_{2}^{\text{(in)}}|_{\tilde{r}^{\ast}}+{B}_{2}\partial_{{r}^{\ast}}h_{2}^{\text{(up)}}\big{|}_{\tilde{r}^{\ast}},$ (42)
${A}_{3}h_{3}^{\text{(in)}}+{B}_{3}h_{3}^{\text{(up)}}\big{|}_{\bar{r}^{\ast}}=\mathcal{A}_{2}h_{2}^{\text{(in)}}\big{|}_{\bar{r}^{\ast}}+{B}_{2}h_{2}^{\text{(up)}}\big{|}_{\bar{r}^{\ast}},$ (43)
${A}_{3}\partial_{r^{\ast}}h_{3}^{\text{(in)}}\big{|}_{\bar{r}^{\ast}}+{B}_{3}\partial_{r^{\ast}}h_{3}^{\text{(up)}}\big{|}_{\bar{r}^{\ast}}$ (44)
@@
Now the coefficients $A_{3}$ and $B_{3}$ can be determined from above matching conditions. To read out $A_{H}^{+}$ and $A_{H}^{-}$ for the asymptotical behavior of the solution at $r^{\ast}\rightarrow-\infty$ in (35) from the solution at $i=3$ in (39), one more step is to rewrite $h_{3}^{\text{(up)}}$ into a combination of the following plane wave forms by employing the identity of the hypergeometric functions
Table 1: The comparison of the fundamental mode frequency with the shooting method (SM) and continued fraction method (CFM). We fixed $m=1$, $\alpha=1/2$, $r_{0}=15$.
$\Omega_{0}=0.1$ $\Omega_{0}=0.2$
$m$ SM CFM SM CFM
-2 $0.3027-0.0055i$     $\,\,\,\,0.3029-0.0055i$   $\,\,0.3231-0.0031i$   $\,\,0.3231-0.0031$
-1 $0.2238-0.0136i$   $\,\,0.2238-0.0137i$   $\,\,0.2392-0.0152i$   $\,\,0.2392-0.0152i$
1 $0.2694-0.0019i$   $\,\,0.2696-0.0020i$   $\,\,0.2914+0.0013i$   $\,\,0.2915+0.0013i$
2 $0.3602-0.0076i$   $\,\,0.3605-0.0075i$   $\,\,0.3906-0.0010i$   $\,\,0.3907-0.0011i$
3 $0.4433-0.0136i$   $\,\,0.4461-0.0147i$   $\,\,0.4804-0.0024i$   $\,\,0.4805-0.0023i$
$\Omega_{0}=0.001$ $\Omega_{0}=0.01$
$m$ SM CFM SM CFM
-2 $0.2379-0.1371i$     $\,\,\,\,0.2362-0.1309i$   $\,\,0.2496-0.0618i$   $\,\,0.2487-0.0683i$
-1 $0.1475-0.1106i$   $\,\,0.1483-0.1068i$   $\,\,0.1694-0.0492i$   $\,\,0.1713-0.0530i$
1 $0.1811-0.1489i$   $\,\,0.1706-0.1292i$   $\,\,0.2058-0.0614i$   $\,\,0.2023-0.0668i$
2 $0.2693-0.1775i$   $\,\,0.2619-0.1684i$   $\,\,0.2903-0.0852i$   $\,\,0.2974-0.0891i$
${}_{2}F_{1}(a,b;c;y)=$ $\frac{\Gamma(c)\Gamma(c-a-b)}{\Gamma(c-a)\Gamma(c-b)}\,_{2}F_{1}(a,b;a+b+1-c;1-y)$
$+\frac{\Gamma(c)\Gamma(a+b-c)}{\Gamma(a)\Gamma(b)}(1-y)^{c-a-b}\,_{2}F_{1}(c-a,c-b;1+c-a-b;1-y),$ (45)
@@
×$\times\,_{2}F_{1}\left(-\beta,\beta+1;1-\mu;\frac{1-y}{2}\right),$ (46)
@@
$h_{3}^{\text{(}up)}(r^{\ast}\rightarrow-\infty)={N}e^{-\mu_{3}r^{\ast}/s_{3}}+{M}e^{\mu_{3}r^{\ast}/s_{3}}$ (47)
@@
${M}=\frac{\Gamma(1-\mu_{3})\Gamma(-\mu_{3})}{\Gamma(1-\mu_{3}+\beta_{3})\Gamma(-\mu_{3}-\beta_{3})}.$ (48)
@@
$A_{H}^{-}={A}_{3}+{N}{B}_{3},$ (49)
$A_{H}^{+}={M}+{B}_{3}.$ (50)
When the mass shell is absent, $A_{3}$ and $B_{3}$ are found in the matching conditions with ${A}_{2}=0$, ${B}_{2}=1$. The spectrum then can be found by $A_{H}^{+}/A_{H}^{-}=0$ Macedo et al. (2018); Torres et al. (2022). Alternatively, one can start from the $i=3$ region and use (32) to determine $A_{3}=1$ and $B_{3}=0$. The matching points are still at $\bar{r}^{\ast}$ and $\tilde{r}^{\ast}$. Using the above identity rewrites $h_{1}^{\text{(}in)}$ at $r^{\ast}\rightarrow\infty$. Then $A^{(+)}_{\infty}$ and $A^{(-)}_{\infty}$ can be read off from (32). The requirement of $A^{(-)}_{\infty}/A^{(+)}_{\infty}=0$ gives the eigenfequency.
Next, we made an extension of the continued fraction method originated from Leaver (1990) for the case of a mass shell around the black hole. We follow the numerical treatments of calculating the QNMs and the Regge pole spectrum for a black hole surrounded by thin-shell matter in Ref. Torres (2023); Torres et al. (2023) The solution of the Klein-Gordon equation (2) can be written as a series expansion around the point $r=b$ selected to be outside the mass shell where $b>r_{0}+1/\alpha$ in (9).
After substituting it into the (2), we have four-term recurrence relation given by
$\alpha_{k}a_{k+1}+\beta_{k}a_{k}+\gamma_{k}a_{k-1}+\delta_{k}a_{k-2}=0,$ (51)
@@
$\alpha_{k}=$ $-\frac{4\left(b^{2}-1\right)^{2}k(k+1)}{b}\,,$ (52)
$\beta_{k}=$ $\frac{8\left(b^{2}-1\right)k\left[-ib^{3}\omega+\left(b^{2}-3\right)k+1\right]}{b},$ (53)
$\gamma_{k}=$ $6b\left[4k(2k-3)+5\right]-\frac{5\left[4k(3k-5)+9\right]}{b}$
$-4b\left(l^{2}+1\right)m^{2}-16ib^{2}(k-1)\omega$
$+b^{3}\left[-4(k-1)k+4m(2l\omega+m)-1\right]\,,$ (54)
$\delta_{k}=$ $\frac{20(3-2k)^{2}}{b}+8ib^{2}(k-2)\omega$
$+4b\left[-8k^{2}+22k+2\left(l^{2}+1\right)m^{2}-15\right].$ (55)
with the initial conditions
$a_{0}=e^{-i\omega r^{\ast}(b)}H_{\omega,m}(b),$ (56)
$a_{1}=be^{-i\omega r^{\ast}(b)}\left[H^{\prime}_{\omega,m}(r)-\frac{i\omega}{r^{2}-1}H_{\omega,m}(r)\right]_{r=b},$ (57)
which can be found numerically by integrating (2) from the horizon $r=1$ up to $r=b>r_{0}+1/\alpha$.
Figure 6: The absolute value of time-domain profile with varying $\Omega_{0}$. The inset figure shows the fundamental QNM frequency with the corresponding value of $\Omega_{0}$. Other parameters are fixed as $m=1$, distance $r_{0}=15$ and $\alpha=1/2$.
Furthermore, one can use Gaussian elimination to reduce (51) to the three-term recurrence relation
$\alpha_{0}^{\prime}a_{1}+\beta_{0}^{\prime}a_{0}=0,$ (58)
$\alpha_{k}^{\prime}a_{k+1}+\beta_{k}^{\prime}a_{k}+\gamma_{k}^{\prime}a_{k-1}=0,\quad k=1,2,\cdots.$ (59)
@@
$\alpha_{k}^{\prime},\,\beta_{k}^{\prime}$ and $\gamma_{k}^{\prime}$ can be expressed in terms of $\alpha_{k},\,\beta_{k},\,\gamma_{k}$ and $\delta_{k}$, given by
$\alpha_{0}=\alpha_{0}^{\prime},\,\beta_{0}^{\prime}=\beta,\,\gamma_{0}^{\prime}=\gamma_{0}$ (60)
$\alpha_{1}=\alpha_{1}^{\prime},\,\beta_{1}^{\prime}=\beta_{1},\,\gamma_{1}^{\prime}=\gamma_{1}$ (61)
$\alpha_{k}^{\prime}=\alpha_{k},\,\beta_{k}^{\prime}=\beta_{k}-\alpha_{k-1}^{\prime}\delta_{k}/\gamma_{k-1}^{\prime},$
and $\gamma_{k}^{\prime}=\gamma_{k}-\beta_{k-1}^{\prime}\delta_{k}/\gamma_{k-1}^{\prime}.$ (62)
Regarding to the issue of convergence of series expansion (S34), we refer the study of Benhar et al. (1999) where mentioned that $a_{k}$ is a minimal solution to the recurrence relation when $b/2<r_{0}+1/\alpha<b$ that in turn give the continued fraction as
$\frac{a_{1}}{a_{0}}=-\frac{\gamma_{1}^{\prime}}{\beta_{1}^{\prime}-}\frac{\alpha_{1}^{\prime}\gamma_{2}^{\prime}}{\beta_{2}^{\prime}-}\frac{\alpha_{2}^{\prime}\gamma_{3}^{\prime}}{\beta_{3}^{\prime}-}\cdots.$ (63)
The complex-frequency roots of the above equation corresponding to different Rabi coupling strengths are shown in Table I, demonstrating that this modified continued fraction method can calculate not only quasibound state but also the destabilized QNMs Cheung et al. (2022).
Appendix C NUMERICALLY SOLVING THE TIME-DEPENDENT EQUATION IN (6) FOR $H_{m}(r,t)$
$H_{\omega,m}(r)=e^{i\omega r^{\ast}(r)}\sum_{k=0}^{\infty}a_{k}\left(1-\frac{b}{r}\right)^{k}.$ (64)